Log X + log(6x+1)=log2

Answer by Boreal(15235)

multiply when logs are added
log x(6x+1)= log 2
raise to 10th power to remove the logs
x(6x+1)=2
6x^2+x-2=0
(3x+2)(2x-1)=0
x=-2/3, 1/2 but logs of negative numbers don't exist
log 1/2 + log 4=log 2
that is log 2^(-1)+ log 2^2=log (2)^1 to check
x=1/2